# It’s Not You, It’s Clipping: A Soft Trust-Region via Probability Smoothing for LLM RL

### v1

@@ 3 Experimental Setup
3.1 Model and Prompt Formatting
We fine-tune the open source causal LMs Qwen2.5-0.5B and -1.5B (Qwen Team, 2024) using their own tokeniser. All runs use identical tokenisation and prompt formatting. Each sample is formatted with a system instruction followed by the user problem text. We use the model’s native chat template via tokenizer.apply_chat_template(…, add_generation_prompt=True) to append the assistant header and ensure the model completes in the assistant role.
@@ 3.1 Model and Prompt Formatting
When decoding completions, we set max_completion_length=128. We do not enforce any additional stop strings beyond the template EOS.
@@
We use this formatting to encourage stepwise reasoning and finish with a single, numeric final answer, which can be more easily extracted for calculating the reward.
3.2 Mathematical Reasoning Datasets
We train on GSM8K (standard train/test split) (Cobbe et al., 2021) and evaluate in-domain on GSM8K and out-of-distribution on ASDiv (Miao et al., 2020), SVAMP (Patel et al., 2021), and MATH-500 (Lightman et al., 2023). These benchmarks span basic arithmetic word problems (ASDiv), robustness to linguistic perturbations (SVAMP), and competition-level reasoning (MATH-500; sampled from MATH (Hendrycks et al., 2021)). Following Minerva (Lewkowycz et al., ) and OpenWebMath (Paster et al., 2024), we restrict evaluation to problems with numeric final answers to enable automatic verification. For GSM8K training, we split the published train set into 7000 train and 472 validation examples.
3.3 Reward Function
@@
Our rewards follow the commonly used correctness-based setup (Lewkowycz et al., ; Paster et al., 2024; DeepSeek-AI et al., 2025): R=1R{=}1 for exact numeric correctness within $10^{-6}$ tolerance, plus a +0.05+0.05 shaping bonus if the output matches the format “#### <<number>>”; values are constrained to $[0,1]$ giving $\{0,0.05,1\}$. We first attempt to extract the number from the requested format; if this is not present, we fall back to the last numeric token in the completion.
3.4 Training
[…]
3.5 Methods
We compare GR-PSPO to two GRPO variants: GRPO-clip (standard clipped ratio) and GRPO-noclip (single iteration over the data). GR-PSPO and GRPO-clip use 2 iterations (data reuse), while GRPO-noclip uses 1 iteration (setting the ratio to 1). To reduce memory and match common defaults, we set $\beta=0$ in the KL penalty (Hugging Face, ) (cf. Eqs. 2, 8). We also compare with two baselines: (i) the base model with the same decoding settings; (ii) SFT on GSM8K using trl’s SFTTrainer with cross-entropy and the identical prompt template.
3.5.1 Hyperparameters
We list some of the hyperparameters in Table  1. We ran hyperparameter tuning for each method using a small grid-search across the learning rate and clipping range/smoothing strength. We ran each method for 500 global training steps on the GSM8K training set, using the evaluation reward to determine the optimal parameter values. We kept the memory-related parameters consistent across all methods, with bf16=True, num_generations=4, per_device_train_batch_size=4, and gradient_accumulation_steps=16222Given that we used 2 GPUs for training, the effective batch size was 64.. We also kept the decoding parameters consistent across all methods, during training: temperature=0.8 and top-p=0.9. All other hyperparameters are the GRPOTrainer defaults(Hugging Face, ), including the AdamW optimiser, the KL coefficient $\beta=0$ and 3 training epochs.
Table 1: Training hyperparameters for the different methods used when fine-tuning Qwen2.5-0.5B and -1.5B on GSM8K training dataset.
Parameter GRPO-noclip GRPO-clip GR-PSPO
Number of Iterations 1 2 2
Qwen2.5-0.5B Parameters
Learning rate $1\times 10^{-6}$ $5\times 10^{-6}$ $5\times 10^{-7}$
Clipping Range ($\varepsilon$) N/A 0.1 N/A
Smoothing Strength ($\alpha$) N/A N/A 0.1
Qwen2.5-1.5B Parameters
Learning rate $1\times 10^{-6}$ $5\times 10^{-7}$ $5\times 10^{-7}$
Clipping Range ($\varepsilon$) N/A 0.2 N/A
Smoothing Strength ($\alpha$) N/A N/A 0.1
3.6 Evaluation
We evaluate each test set across temperatures $T\in\{0.0,0.2,0.4,0.6,0.8\}$ with top-p=1.0p{=}1.0. We report zero-shot Top-1 accuracy with 95% confidence intervals. We use Top-1 at $T=0$ as it reflects single-answer deployment settings and is deterministic.
We also assess response quality by scoring 5 metrics on a scale of 151-5 using an LLM-as-Judge (validating results by sampling a subset of the responses ourselves). The metrics scored are: overall quality; constraint adherence (format fidelity, steps present); logical coherence (no contradictions, consistent rationale); mathematical soundness (valid operations/derivations); and clarity (concise, well-structured). Prompts are in App. B.
4 Results
At T=0T{=}0 (greedy, deterministic), GR-PSPO improves Top-1 over GRPO-clip by +22.1+22.1pp on GSM8K for the 0.5B model (39.7 vs. 17.6) and +21.6+21.6pp for the 1.5B model (59.4 vs. 37.8), while remaining comparable to GRPO-noclip. There are positive gains on ASDiv/SVAMP (approx. 20pp on 0.5B and 7pp and 12pp on 1.5B); MATH-500 shows minimal improvements.
Tables 2 and  3 report our full results alongside published performances on similar model sizes; we omit published results which are below $15\%$ accuracy from the table (Brown et al., ; Ho et al., 2023; Zhuang et al., 2025; Ho et al., 2023) for visual clarity. The full table can be found in the Appendix A.
0.5B.
Results on the 0.5B model are shown in Table 2. GR-PSPO and GRPO-noclip demonstrate similar performance across all datasets ($\leq$1pp difference; CIs overlap), while GR-PSPO is far ahead of GRPO-clip across datasets (e.g., GSM8K $39.7$ vs. $17.6$). On MATH-500, improvements vs. clipping are smaller (16.8 vs. 10.3). GR-PSPO and GRPO-noclip both outperform SFT on all datasets, but SFT does outperform GRPO-clip. Compared to literature at similar scale Luo et al. (2025) outperforms our methods on both GSM8K and MATH, using a slightly larger model (0.70.7B vs. 0.50.5B). We note that our models were only fine-tuned on GSM8K, unlike Luo et al. (2025), which fine-tunes on both GSM8K and MATH. Zhuang et al. (2025) performs the best on MATH-500, with $23.6\%$.
1.5B.
Results on the 1.5B model are shown in Table 3. GR-PSPO attains the best T=0T{=}0 Top-1 on GSM8K (59.4) and MATH-500 (25.6), with GRPO-noclip slightly ahead on ASDiv/SVAMP; both outperforming GRPO-clip. Confidence intervals are larger at this scale, effectively covering the gap between GR-PSPO and GRPO-noclip. Compared with the published results, Luo et al. (2025) is the closest result, with slightly poorer performance on both datasets.
Model Size Fine-tuning Dataset GSM8K ASDiv SVAMP MATH-500
WizardMath-GPT2-Small (Luo et al., 2025) 0.1B GSM8K & MATH 26.4 - - 12.3*
WizardMath-GPT2-Medium (Luo et al., 2025) 0.3B GSM8K & MATH 38.7 - - 15.6*
Qwen2.5-0.5B+RL (Zhuang et al., 2025) 0.5B GSM8K 34.1 - - 23.6
WizardMath-GPT2-Large (Luo et al., 2025) 0.7B GSM8K & MATH 50.1 - - 21.2*
FlanT5-Large + Specialised (Fu et al., ) 0.76B GSM8K 20.2 23.8 20.4 -
GPT-2 (Socratic CoT) (Shridhar et al., 2023) 0.77B GSM8K + Socratic CoT 21.1 - - -
GPT-2 (Socratic CoT) (Shridhar et al., 2023) 0.77B SVAMP + Socratic CoT 21.1 - 23.6 -
Our Results
Qwen2.5-0.5B 0.5B - 8.3 50.8 26.3 10.2
SFT 0.5B GSM8K 28.6 57.6 38.1 14.2
GRPO-clip 0.5B GSM8K 17.6 42.8 34.3 10.3
GRPO-noclip 0.5B GSM8K 40.7 68.3 53.9 17.2
GR-PSPO 0.5B GSM8K 39.7 68.1 53.2 16.8
Table 2: Performance of similar-size smaller language models (<1<1B parameter) from the literature, compared with our results fine-tuning Qwen2.5-0.5B. We report our model accuracy from the zero-shot Top-1 accuracy at temperature $T=0$.
*These models were evaluated against MATH rather than MATH-500. Where MATH-500 is a subset of MATH, we still compare the results here.
Model Size Fine-tuning Dataset GSM8K ASDiv SVAMP MATH-500
GPT-2-XL (Brown et al., ) 1.5B GSM8K & MATH 15.4 - - 6.9*
WizardMath-GPT2-XL (Luo et al., 2025) 1.5B GSM8K & MATH 58.9 - - 25.4*
Our Results
Qwen2.5-1.5B 1.5B - 1.6 2.4 1.7 5.1
GRPO-clip 1.5B GSM8K 37.8 70.9 58.4 14.9
GRPO-noclip 1.5B GSM8K 57.9 80.4 74.9 25.2
GR-PSPO 1.5B GSM8K 59.4 77.7 70.3 25.6
Table 3: Performance of similar-size large language models (>1>1B parameter) from the literature, compared with our results fine-tuning Qwen2.5-1.5B. We report our model accuracy from the zero-shot Top-1 accuracy at temperature $T=0$.
*These models were evaluated against MATH rather than MATH-500. Where MATH-500 is a subset of MATH, we still compare the results here.
Response Quality:
LLM-as-Judge scores (1–5) favour GR-PSPO’s responses on GSM8K/ASDiv/MATH-500 across overall quality, constraint adherence, logical coherence, mathematical soundness, and clarity; GRPO-noclip is best on SVAMP.
An example of each of our Qwen2.5-0.5B fine-tuned models responses are shown below for a question from the ASDiv dataset:
Prompt:
The base of the parallelogram is 11 inches and its area is 132 Sq. inches. Find the height of the parallelogram.
GRPO-clip:
The height of the parallelogram can be calculated using the formula: height = area/baseheight = 132/11 = 12 inchesThe answer: 12 inches.The answer is: 12 inches reibungsystem You are a careful math solver. Think through the solution and show the steps. Use English only. End the response with the final answer only in the format: ’#### <<final numeric answer only>>
GRPO-noclip:
The height can be found by dividing the area by the base, which gives us 132/11 = 12 inches. #### 12 #### No answer provided and no explanation given. Please calculate and respond with the result. #### No answer provided …
GR-PSPO:
The area of a parallelogram can be calculated as the product of its base and height. Let’s call the base 11 inches and the area 132 square inches. So we have 11 * h = 132. Solving for h, we get h = 132 / 11 = 12 inches. #### 12
Both GRPO-clip and GRPO-noclip have a tendency to leak system instructions, rather than finishing the response early. GRPO-noclip sometimes leaks system instruction-like data, but worded differently than how we prompted the model, this may be from the base model training. Additionally, GRPO-noclip tends to append different alphabet characters, likely due to confusion within the model as the base model is trained on multiple languages.
GR-PSPO tends to produce clearer, shorter responses. It does on occasion duplicate the final value at the end of the message, or append additional ’####’.
Dataset Model Overall Constraints Logic Maths Clarity
GSM8K Qwen2.5-0.5B 2.450 2.650 2.400 2.400 2.350
GSM8K GRPO-clip 2.072 2.486 2.133 2.082 2.009
GSM8K GRPO-noclip 3.579 4.191 3.393 3.302 3.839
GSM8K GR-PSPO 3.720 4.310 3.590 3.520 4.080
ASDiv Qwen2.5-0.5B 3.000 3.000 3.000 3.000 2.950
ASDiv GRPO-clip 2.925 3.218 2.946 2.976 2.479
ASDiv GRPO-noclip 4.380 4.610 4.300 4.330 4.380
ASDiv GR-PSPO 4.490 4.630 4.400 4.450 4.600
SVAMP Qwen2.5-0.5B 1.500 2.600 1.400 1.550 1.450
SVAMP GRPO-clip 2.700 3.340 2.640 2.750 2.420
SVAMP GRPO-noclip 3.356 4.031 3.102 3.226 3.683
SVAMP GR-PSPO 3.260 3.820 2.955 3.106 3.627
MATH-500 Qwen2.5-0.5B 4.053 4.579 4.053 4.000 3.842
MATH-500 GRPO-clip 3.967 4.262 4.010 3.967 3.793
MATH-500 GRPO-noclip 4.222 4.559 4.139 4.199 4.155
MATH-500 GR-PSPO 4.536 4.840 4.493 4.483 4.480
Table 4: Averaged scores of the model responses as judged by the llama-3.3-70b-versatile LLM. Constraints - how well the response followed the task constraints; Logic - how coherent and contradiction-free the logical reasoning was; Maths - how sound and valid the mathematical steps were; Clarity - how concise and clear the response was.
5 Discussion
Compared with ratio clipping and explicit KL regularisation, PSPO provides stability without needing to truncate the surrogate objective, and it does so without adding compute or an extra optimisation objective.
The empirical results demonstrate that GR-PSPO consistently outperforms GRPO-clip on all datasets across both model sizes. Compared to GRPO-noclip, we achieve similar performances on the 0.5B model, within $1\%$ difference, which is within the confidence intervals of each model. Similarly, on the 1.5B model, we perform slightly better on GSM8K and MATH-500, again, all differences are within the confidence interval bands of the models.
Importantly, we notice a distinct difference in response styles, with our LLM-judge ranking GR-PSPO as better in all categories on the GSM8K, ASDiv and MATH-500 datasets. On SVAMP, all methods perform worse, with GRPO-noclip performing best of the bunch. Response quality is important in LLM training for extractability, format fidelity, and tool use. GR-PSPO’s behaviour-anchored smoothing reduces instruction leakage and verbosity while improving clarity and constraint adherence.
Furthermore, Zheng et al. (2025); NormalUhr (2025) note that larger or sparser models (e.g. Mixture-of-Experts), often require larger rollout batch sizes, and to improve sample efficiency mini-batches are necessary. In such cases, GRPO-noclip would be undesirable, as iterating over mini-batches without clipping (or an alternative) would lead to large, unstable steps. Future work to extend empirical testing to larger models would be beneficial to find the limit of GRPO-noclip’s usefulness, and confirm that GR-PSPO continues to demonstrate stable training.
6 Limitations
While our results demonstrate the effectiveness of GR-PSPO on mathematical reasoning tasks, there are limitations in our experimental approach.
Our experimental evaluation is only on mathematical reasoning, where there are binary, objective reward signals. The effectiveness of probability smoothing in domains with more subjective or continuous rewards remains unexplored, and should be considered in future work. This will also provide more insight into the sensitivity of $\alpha$ across domains.
The scale of our experiments is limited to models under 2B parameters. In practice, larger models are normally deployed. Although we demonstrate GR-PSPO performs across two model sizes, future work should consider larger model sizes, as well as different architectures and tokeniser uses.
Finally, GR-PSPO achieves similar quantitative performance to GRPO-noclip, and we have noted that literature(Zheng et al., 2025; NormalUhr, 2025) both have claimed GRPO struggles with larger models and more complex architectures. Empirically comparing GR-PSPO against GRPO in these settings will allow a fuller characterisation of where our method provides practical advantages.

### v2

@@ 3 Experimental Setup
Model and Prompt Formatting We fine-tune the open source causal LMs Qwen2.5-0.5B and -1.5B (Qwen Team, 2024) and also the Qwen2-Math-1.5B model which has specialised domain knowledge of mathematics (Yang et al., 2024) using their own tokeniser. The primary difference between the Math and the base models is that the Math model is a Qwen2-1.5B model which is then pretrained on a mathematics-specific corpus using supervised fine-tuning.
We use these three models as they are openly available, allow us to test scale across different, smaller sizes and enables us to compare models with and without existing mathematics knowledge. All runs use identical tokenisation and prompt formatting. Each sample is formatted with a system instruction followed by the user problem text. We use the model’s native chat template via tokenizer.apply_chat_template(…, add_generation_prompt=True) to append the assistant header and ensure the model completes in the assistant role, consistent with how the models are trained to be used.
When decoding completions, we set max_completion_length=128 with a maximum prompt length of 512 tokens, giving 640 tokens overall. Token count in LLM fine-tuning is a big contributing factor on the compute required (approx. 4 times the token count memory required for training), so we chose 128 completion tokens to balance computational constraints whilst allowing room for the response to include reasoning steps. This is also the approximate length of the GSM8K gold answers. We do not enforce any additional stop strings beyond the template defined end of string token.
Training All methods are trained using the same effective batch size and decoding settings. We train each method across 3 seeds, for the same number of tokens for fair comparison of the different methods; the best-validation checkpoint model is used for evaluations.
For the base models (Qwen2.5-0.5B and -1.5B), which lack specialised knowledge, we trained for 150M tokens. For Qwen2-Math-1.5B, which has existing mathematical knowledge to build on, we train for 75M tokens. Models with existing domain knowledge typically require less training to refine.
[…]
Table 1: Trust region methods evaluated. All use GRPO’s group relative advantages.
Method Trust Region Mechanism Implementation
GRPO-clip IS ratio clipping TRL baseline
[…]
P3O Sigmoidal soft clipping (SCOPIC) Adapted to GRPO
TOPR Asymmetric advantage gating Adapted to GRPO
SAPO Asymmetric sigmoidal scaling Backported from TRL
GR-PSPO Probability smoothing Novel (ours)
Hyperparameter Tuning To tune the hyperparameters, we used a grid-search across learning rate, clipping range, smoothing alpha strength, and tau values. Memory-related and decoding parameters were fixed across all methods. More details on the tuning method and the parameters used for each method can be found in Appendix A.
Evaluation We evaluate each test set across temperatures $T\in\{0.0,0.2,0.4,0.6,0.8\}$ with top-p=1.0p{=}1.0. We report the mean zero-shot Top-1 accuracy with 95% confidence intervals. We use Top-1 at $T=0$ as it reflects single-answer deployment settings and is deterministic. We calculate the accuracy using the true accuracy of a response, i.e. if anywhere in the response the correct answer appears. We also evaluate the reward function style accuracy as well.
3.1 Mathematical Reasoning Datasets
The base models were fine-tuned on GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) (using the standard train/test split) and evaluated in-domain on the test sets and out-of-distribution on ASDiv (Miao et al., 2020) and SVAMP (Patel et al., 2021). These benchmarks span basic arithmetic word problems (ASDiv), robustness to linguistic perturbations (SVAMP), and competition-level reasoning (MATH; sampled from MATH (Hendrycks et al., 2021)). Following Minerva (Lewkowycz et al., 2022) and OpenWebMath (Paster et al., 2024), we restrict evaluation to problems with numeric final answers to enable automatic verification. For GSM8K training, we split the published training set into 7000 train and 472 validation examples.
3.1.1 System Prompt
@@
We use this prompting to encourage stepwise reasoning and finish with a single, numeric final answer, which can be more easily extracted for calculating the reward.
3.1.2 Reward Function
@@
Our rewards follow the commonly used correctness-based setup (Lewkowycz et al., 2022; Paster et al., 2024; DeepSeek-AI et al., 2025): R=1R{=}1 for exact numeric correctness within $10^{-6}$ tolerance, plus a +0.05+0.05 shaping bonus if the output matches the format “#### <<number>>”; values are constrained to $[0,1]$ giving $\{0,0.05,1\}$. We first attempt to extract the number from the requested format; if this is not present, we fall back to the last numeric token in the completion.
Table 2: Comparison of our results across all the model sizes on mathematical reasoning benchmarks. The accuracy presented is the true accuracy (i.e. has the correct answer anywhere in the response, even if it provides an incorrect answer at the end). Accuracy reported as average over 3 seeds, evaluated in greedy setting with T=0.0, and top_p=1.0.
Model Size Training Accuracy (%)
Data GSM8K ASDiv SVAMP MATH
Qwen2.5-0.5B 0.5B N/A 16.9 42.7 42.0 30.2
Qwen2.5-0.5B+GRPO-clip 0.5B GSM8K 54.1 77.8 66.6 43.0
Qwen2.5-0.5B+GRPO-noclip 0.5B GSM8K 56.0 77.9 70.0 42.6
Qwen2.5-0.5B+P3OChen et al. (2022) 0.5B GSM8K 44.5 68.5 58.3 38.9
Qwen2.5-0.5B+SAPOGao et al. (2025) 0.5B GSM8K 31.8 62.0 51.8 36.2
Qwen2.5-0.5B+GR-PSPO 0.5B GSM8K 53.0 78.8 69.7 41.9
Qwen2.5-0.5B+GRPO-clip 0.5B GSM8K+MATH 49.4 76.5 62.1 34.4
Qwen2.5-0.5B+GRPO-noclip 0.5B GSM8K+MATH 54.1 76.6 56.8 36.2
Qwen2.5-0.5B+P3OChen et al. (2022) 0.5B GSM8K+MATH 50.1 75.0 62.8 36.1
Qwen2.5-0.5B+SAPOGao et al. (2025) 0.5B GSM8K+MATH 50.3 76.6 63.3 34.5
Qwen2.5-0.5B+TOPRRoux and Bellemare (2025) 0.5B GSM8K+MATH 2.8 3.2 7.3 8.6
Qwen2.5-0.5B+GR-PSPO 0.5B GSM8K+MATH 50.9 75.9 65.7 37.5
Qwen2.5-1.5B 1.5B N/A 5.2 5.0 4.0 17.6
Qwen2.5-1.5B+GRPO-clip 1.5B GSM8K+MATH 73.7 88.8 83.3 45.9
Qwen2.5-1.5B+GRPO-noclip 1.5B GSM8K+MATH 75.4 87.9 83.7 48.0
Qwen2.5-1.5B+GR-PSPO 1.5B GSM8K+MATH 73.8 89.4 82.7 50.0
Qwen2-Math-1.5B 1.5B N/A 24.7 61.8 56.3 34.8
Qwen2-Math-1.5B+GRPO-noclip 1.5B GSM8K+MATH 78.4 92.8 84.6 58.3
Qwen2-Math-1.5B+GRPO-clip 1.5B GSM8K+MATH 70.3 91.6 82.0 52.3
Qwen2-Math-1.5B+P3OChen et al. (2023) 1.5B GSM8K+MATH 78.8 93.0 84.7 55.8
Qwen2-Math-1.5B+SAPOGao et al. (2025) 1.5B GSM8K+MATH 79.2 93.8 85.3 59.3
Qwen2-Math-1.5B+GR-PSPO 1.5B GSM8K+MATH 79.9 94.0 85.7 59.6
4 Performance of PSPO compared to the state-of-the-art
We evaluate GR-PSPO against multiple trust region baselines across mathematical reasoning tasks. Table 2 shows results across all model configurations. Our main finding is that GR-PSPO’s advantages grow with model size and domain specialisation, validating our hypothesis that probability smoothing is most effective when refining models that already have substantial knowledge, when stability is most important.
4.1 Model Capability and Trust Region Performance
GR-PSPO’s advantage scales with prior domain knowledge. On Qwen2.5-0.5B, differences between methods are small. On Qwen2.5-1.5B, GR-PSPO achieves the strongest MATH performance (+4.1pp over GRPO-clip). On Qwen2-Math-1.5B, a model with existing mathematical domain knowledge, GR-PSPO outperforms on all four benchmarks.
This scaling is reflected in GSM8K performance, where GR-PSPO transitions from -$\text{-}3.2$% relative to the best per dataset (0.5B) to -$\text{-}1.6$% (1.5B) to +0.7% (Math-1.5B).
Math-Specialised Model
On Qwen2-Math-1.5B, GR-PSPO achieves the strongest performance across all benchmarks. At $T=0$, GR-PSPO attains 79.9% on GSM8K (versus 78.4% for GRPO-noclip), 94.0% on ASDiv, 85.7% on SVAMP, and 59.6% on MATH.
The consistency across tasks—from arithmetic word problems (ASDiv, SVAMP) to competition-level reasoning (MATH)—suggests that probability smoothing generalises across difficulty levels when refining models with a base competence.
4.2 Base Models
On base Qwen2.5 models without prior mathematical training, GR-PSPO’s advantages are most pronounced on MATH, the most challenging benchmark. For Qwen2.5-0.5B trained on GSM8K+MATH, GR-PSPO achieves the highest MATH accuracy (37.5% vs. 36.2% for GRPO-noclip) but trails on GSM8K (50.9% vs. 54.1%). For Qwen2.5-1.5B, GR-PSPO achieves 50.0% on MATH versus 45.9% for GRPO-clip (+4.1pp), with comparable GSM8K performance.
When fine-tuning a base model on a new task, there is less existing knowledge to preserve, and aggressive updates may accelerate learning. When refining a model that already has domain expertise, stability becomes critical. GR-PSPO’s behaviour-anchored smoothing provides this property, explaining its consistent gains on Qwen2-Math.
4.3 Comparison to alternative methods.
GR-PSPO generally outperforms P3O and SAPO on both GSM8K and MATH across model sizes (Table 2). On Math-1.5B, GR-PSPO achieves 59.6% on MATH versus 55.8% for P3O (+3.8pp). We also adapted TOPR (Roux and Bellemare, 2025) to the GRPO framework, but the model seems to collapse during training, leading to below base model performance. TOPR was originally designed for REINFORCE, so the model may of collapsed due to the group-relative advantage formulation in GRPO.
4.4 Comparison to published results.
Our results outperform prior work at similar scales (Table 3). Luo et al. (2025) report 50.1% GSM8K and 21.2% MATH on WizardMath-Large (0.7B); our 0.5B GR-PSPO achieves 50.9% and 37.5% respectively. Our Qwen2-Math-1.5B results (79.9% GSM8K, 59.6% MATH) substantially exceed WizardMath-GPT2-XL (58.9%, 25.4%) at comparable scale. More notably, our Qwen2-Math-1.5B results outperform the Llama 3 8B model from Roux and Bellemare (2025), achieving +4.5pp on GSM8K and +36.9pp on MATH despite being substantially smaller. Direct comparison is complicated by differences in evaluation methodology; notably, our context window is much shorter (640 tokens with 128 completion tokens versus 2000+ tokens in prior work, or 128 completion tokens vs. the 512 completion tokens). We evaluate on the harshest settings, and still have competitive performance.
Table 3: Comparison on mathematical reasoning benchmarks. Reward Source: Rule = rule-based correctness function, Model = trained reward model. Reward Type: O = outcome (final answer), P = process (intermediate reasoning steps). The accuracy presented is the true accuracy (i.e. has the correct answer anywhere in the response, even if it provides an incorrect answer at the end). Note: “—” indicates data could not be found; “N/A” indicates not applicable.
Model Size Training Reward Evaluation Accuracy (%)
Data Method Source Type Prompting Decoding Tokens GSM8K MATH
Smaller Language Models << 1B parameters.
WizardMath-Small (Luo et al., 2025) 0.1B GSM8K+MATH PPO Model P+O - T=0 2048 26.4 12.3
WizardMath-Medium (Luo et al., 2025) 0.3B GSM8K+MATH PPO Model P+O - T=0 2048 38.7 15.6
WizardMath-Large (Luo et al., 2025) 0.7B GSM8K+MATH PPO Model P+O - T=0 2048 50.1 21.2
Qwen2.5-0.5B+RL (Zhuang et al., 2025) 0.5B GSM8K GRPO Rule O - T=0.6, top_p=0.95 32768 34.1 23.6
FlanT5-L Specialized (Fu et al., 2023) 0.76B GSM8K Distill N/A N/A Few-shot - 8192 20.2 -
GPT-2 (Socratic) (Shridhar et al., 2023) 0.77B GSM8K+SCoT Distill N/A N/A Socratic CoT - - 21.1 -
Qwen2.5-0.5B+GRPO-noclip 0.5B GSM8K+MATH GRPO (iterations=1) Rule O Zero-shot T=0, top_p=1.0 640 54.1 36.2
Qwen2.5-0.5B+GR-PSPO 0.5B GSM8K+MATH GR-PSPO Rule O Zero-shot T=0, top_p=1.0 640 50.9 37.5
Mid-to-Large Language Models >> 1B parameters.
GPT-2-XL (Brown et al., 2020) 1.5B GSM8K+MATH N/A N/A N/A Few-Shot - 2048 15.4 6.9
WizardMath-GPT2-XL (Luo et al., 2025) 1.5B GSM8K+MATH PPO Model P+O - T=0 2048 58.9 25.4
Llama 3 8B+TOPR Roux and Bellemare (2025) 8B GSM8K+MATH TOPR Rule O Few-Shot - 512 (completion tokens) 75.4 22.7
Qwen2.5-1.5B+GRPO-noclip 1.5B GSM8K+MATH GRPO (iterations=1) Rule O Zero-shot T=0, top_p=1.0 640 75.4 48.0
Qwen2.5-1.5B+GR-PSPO 1.5B GSM8K+MATH GR-PSPO Rule O Zero-shot T=0, top_p=1.0 640 73.8 50.0
Qwen2-Math-1.5B+GR-PSPO 1.5B GSM8K+MATH GR-PSPO Rule O Zero-shot T=0, top_p=1.0 640 79.9 59.6
Smoothing Alpha Sensitivity
We ablate $\alpha$ on Qwen2-Math-1.5B (Appendix B). Performance peaks around $\alpha$ = 0.4.
Preliminary larger model results (7B).
To further consider scaling behaviour we ran one exploratory seed for a reduced number of tokens on 7B models. The Qwen2.5-7B run shows a similar trend as our smaller-scale experiments, with: GR-PSPO outperforming on MATH. For example, PSPO’s top-1 (true) accuracy is 55.8% on MATH compared with the base model’s 33.9%, and GRPO-clip’s 53.6%. GRPO-noclip has an accuracy of 0% indicating complete collapse. These 7B measurements are single-seed and exploratory, so we do not use them to change conclusions drawn from the 3-seed experiments.
5 Discussion and Limitations
Unlike clipping which truncates the surrogate objective and discards gradient information outside the allowed range, PSPO contracts ratios toward 1 through probability smoothing while preserving non-zero gradients everywhere. This is achieved without additional compute, memory overhead, or extra optimisation objectives, the only change is a linear transformation of the importance ratio.
Our results across three model configurations validate that probability smoothing is most effective when stability is required, particularly when refining a model with existing domain knowledge.
This is shown by performance improvements scaling with domain specialisation. On general base models (Qwen2.5-0.5B/1.5B), GR-PSPO shows advantages primarily on the hardest benchmark (MATH). On the specialised model (Qwen2-Math-1.5B), GR-PSPO outperforms consistently across all tasks, suggesting that PSPO’s benefits are clearest when there is existing knowledge to preserve.
Additionally, performance improves with scale. The progression from 0.5B (mixed results) to 1.5B (strong MATH performance) to Math-1.5B (consistent gains) demonstrates that PSPO’s advantages compound as model size grows. On MATH, GR-PSPO transitions from -1.1% compared to clipping (0.5B) to +4.1pp absolute advantage (1.5B).
When fine-tuning very small base models (0.5B), the cost of discarding gradient information may be less impactful. But as models grow and accumulate knowledge, each sample becomes more costly to generate, and preserving learning signal everywhere becomes critical. PSPO’s non-zero gradients (Proposition 2.5) provide exactly this property. Unlike clipping’s hard boundary, PSPO’s behaviour-anchored smoothing provides a soft constraint that scales naturally with how far the policy deviates. Sigmoid-based alternatives such as P3O and SAPO are more vulnerable to saturation at extreme ratios and asymmetric boundaries require additional hyperparameters (SAPO and TOPR each require at least two $\tau$ values; TOPR can require up to four).
Compared to GRPO-noclip, GR-PSPO achieves similar or better performance while enabling multi-epoch training. GRPO-noclip avoids the need for trust regions by using only a single iteration over each batch, ensuring the importance ratio is always 1. This can work well when fresh on-policy samples are cheap, but generating rollouts from LLMs is computationally expensive. When reusing data, the policy diverges from the behaviour policy, making trust region mechanisms essential for stability. Additionally, Zheng et al. (2025) and NormalUhr (2025) note that larger models often require processing rollout batches in mini-batches to fit in memory, making single-iteration training impractical. PSPO enables stable multi-epoch training without sacrificing gradient information.
Refining a base model with domain knowledge is one of the most relevant for production LLM deployment. In practice, organisations fine-tune capable foundation models that already possess substantial knowledge and capabilities. The goal is improving performance on specific tasks while preserving the model’s general competence, which is exactly where our results show PSPO excels.